# Segment-Aware Conditioning for Training-Free Intra-Utterance Emotion and Duration Control in Text-to-Speech

### v1

@@ E.1 Baseline and Comparative models
Baseline. IndexTTS2333https://github.com/index-tts/index-tts Zhou et al. (2025) is an autoregressive zero-shot TTS model that supports utterance-level control of emotion and speech duration while maintaining high speech naturalness. It disentangles speaker identity from emotional expression, enabling faithful reconstruction of target timbre and accurate reproduction of the specified emotional style. By incorporating GPT-based latent representations, the model further improves semantic consistency and stability under expressive conditions.
@@
  • CosyVoice2777https://github.com/FunAudioLLM/CosyVoice?tab=readme-ov-file Du et al. (2024b) is an autoregressive TTS model that combines a language model for semantic and prosodic modeling with flow matching for speaker identity reconstruction, utilizing a supervised speech tokenizer to achieve disentangled generation. Notably, it demonstrates superior performance in Chinese compared to English due to its training data distribution.
@@ 36    $\mathcal{M}_{i}[\mathrm{off}+r,\ 0:M\times L_{C}]\leftarrow$-$\infty$
38    $\mathcal{M}_{i}[\mathrm{off}+r,\ L_{C}\times(\mathrm{seg}_{\mathbf{s}}[t]-1):L_{C}\times\mathrm{seg}_{\mathbf{s}}[t]]\leftarrow 0$

### v2

@@ E.1 Baseline and Comparative models
Baseline. IndexTTS2555https://github.com/index-tts/index-tts Zhou et al. (2025) is an autoregressive zero-shot TTS model that supports utterance-level control of emotion and speech duration while maintaining high speech naturalness. It disentangles speaker identity from emotional expression, enabling faithful reconstruction of target timbre and accurate reproduction of the specified emotional style. By incorporating GPT-based latent representations, the model further improves semantic consistency and stability under expressive conditions.
@@
  • CosyVoice2999https://github.com/FunAudioLLM/CosyVoice?tab=readme-ov-file Du et al. (2024b) is an autoregressive TTS model that combines a language model for semantic and prosodic modeling with flow matching for speaker identity reconstruction, utilizing a supervised speech tokenizer to achieve disentangled generation. Notably, it demonstrates superior performance in Chinese compared to English due to its training data distribution.
@@ 36    $\mathcal{M}_{i}[\mathrm{off}+r,\ 0:M\times L_{C}]\leftarrow$-$\infty$
38    $\mathcal{M}_{i}[\mathrm{off}+r,\ L_{C}\times(\mathrm{seg}_{\mathbf{s}}[r]-1):L_{C}\times\mathrm{seg}_{\mathbf{s}}[r]]\leftarrow 0$